OpenStudy (anonymous):

An occupant of a car can survive a crash if the deceleration during the crash is less than 30 g. Calculate the force on a 64 kg person decelerating at this rate.

OpenStudy (anonymous):

It has been awhile so hopefully I am not overlooking something. \[F=ma\] meaning Force = (mass) * (acceleration) Force = Newtons (N), mass = kilograms (kg), acceleration = meters/second^2 You are given the mass which is 64kg. You are given the acceleration as 30 g which is another way of saying 30*g g is the acceleration due to gravity on Earth, it is a constant and it = 9.8m/s^2 Therefore you have the acceleration as well which is 30*9.8=294m/s^2 F = ma so F = (64kg)*(294m/s^2) = 18816 N